OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x2 + 28x - 8 A. Minimum Value: 7 Range y > 7 B. Minimum Value: -106 Range y > -7 C. Minimum Value: -106 Range y > -106 D. Minimum Value: -7 Range y > -7 If you don't mind... @eliassaab do you know how to solve this as well?

Parth (parthkohli):

Can I help?

OpenStudy (anonymous):

Yes please

Parth (parthkohli):

Sir Elias, please continue! I may suggest other methods along the way.

OpenStudy (anonymous):

Please get used to using wolframalpha which will answer most of your questions

OpenStudy (anonymous):

@ParthKohli go ahead a help her

Parth (parthkohli):

There is another way to find the maximum/minimum of the function: find the vertex.

Parth (parthkohli):

@eliassaab Thank you!

OpenStudy (anonymous):

Thank you @eliassaab and... @ParthKohli I'd like to know what that way is if you don't mind...

Parth (parthkohli):

Do you know what a vertex of a graph is? If you don't, I might skip that way.

OpenStudy (anonymous):

I think it's the highest or lowest point on the y-axis... I'm not sure >-<

Parth (parthkohli):

Yes, that's correct! That's the lowest point on a curve.

OpenStudy (anonymous):

Oooh I see... is there a curve from this equation if it's graphed or something?

Parth (parthkohli):

Well, there is a formula for a vertex in the case of a parabola.

Parth (parthkohli):

The vertex is \(\left( -\dfrac{b}{2a}, f\left(\dfrac{-b}{2a}\right)\right)\)

Parth (parthkohli):

OK, never mind that. All hail the Wolfram Alpha Overlords!

OpenStudy (anonymous):

o.o lol what just happened